# An Empirical Bayes Approach for Estimating Skill Models for Professional Darts Players

### v2

@@ 7 Visualizing the Fitted DM Models
In Section 6 we saw that the DM models outperform the benchmark, i.e. non-DM, models in terms of proper scoring rules. In this section we will demonstrate this visually by comparing the BSPN model with an analogous model that is fitted using the raw data. In particular, recall that the BSPN model uses a separate skill model (5) for each double and treble region and fits each of these models using the pseudo-fractions $(\widehat{\alpha}_{k}+x_{jk})/\sum_{l=1}^{K}(\widehat{\alpha}_{l}+x_{jl})$. We will compare the BSPN model with the same model but with all the $\widehat{\alpha}_{k}$’s set to zero so that only the raw fractions are used. We will refer to this latter model as the raw-normal (RN) model.
@@
Severe problems can occur when fitting the RN model to double regions, including even D20, which accounts for over 25% of all doubles in the data-set. CEs for D20 and D14 are displayed in Figure 3. In the case of D20, for example, while all 4 players have observations in D20, S20, and M, these are the only regions covered by Cross, while Cullen (S1), van Gerwen (D5), and Gurney (D5) only have a coverage of 4 when targeting D20. With such limited coverage the likelihood function (which the EM algorithm seeks to optimize) will be quite flat in large neighborhoods and terminating at a reasonable (local) maximum is quite unlikely although it can happen. For example, the RN CEs for van Gerwen and Gurney on D20 look reasonable despite only having a coverage of 4. The quality of fits deteriorate as we consider doubles with even less data. There were only 312 darts targeted at D14, for example, which averages out to approximately 19.5 darts per player. It is no surprise then that the D14 coverages are low for all players including Cross (4), Cullen (3), van Gerwen (4), and Gurney (3) whose fits are displayed in the lower row of Figure 3. Matters are even worse for the odd doubles with D19 (the best of them) having only 154 darts targeting it for an average of 9.6 darts per player and all but 1 player having a coverage of just 2 or 3.
@@ 7.1 An Aside on Correlation Identification
The superior fits provided by the BSPN models bring to light an interesting issue. We see from Figures 3 and 4 that the main axis of the CEs (for the BSPN fits) tends to be aligned with the principal direction of the target region in question. For example, in the case of D14 we see the CEs tend to be oriented in a north-south direction which is also the general orientation of the D14 bed. Similarly, we see the CEs for D20 are oriented east-west which is also the orientation for the D20 bed. Indeed, as may be seen from Figures 10-14 in Appendix D.3, this behavior persists more generally. This is a shortcoming of the bivariate-normal skill model as it stretches the variance along the main axis of the target area in order to fit the fractions of darts that landed in the neighboring doubles / trebles regions. A bivariate distribution with fatter tails, e.g. the bivariate $t$ distribution, might resolve this problem but owing to space considerations as well as other issues related to correlation identification (see below) and interpreting ${\bm{\mu}}$ (see Appendix C.1) we leave this for future research.
Because of this tendency to stretch the variance along the main axis of the target area, we believe the resulting (non-zero) correlations are spurious. This is also related to a problem intrinsic to the data-set. In particular, we do not know the exact $(x,y)$-coordinates of where a dart lands. Instead we only know the region, e.g. S20, T20, etc., and this means identifying the correlation parameter is very difficult. To see this, consider Figure 5 which displays three synthetic data-sets each consisting of 100 darts aimed at T20. It is clear the skill-models generating these data-sets have very different correlations with the first, second, and third data-sets displaying positive, negative, and (approximately) zero correlations, respectively. However, all three data-sets have identical outcomes, i.e. 37, 54, 6, 1, and 2 darts hitting T20, S20, S5, T1, and S1, respectively. As such, the three data-sets would have identical likelihood functions and there is no way for us to identify the correct sign (never mind value) of the correlation. While we have no darts hitting T5 in these artificial data-sets, it should be clear that not having full coverage [^16] It is true, however, that in the limit of infinite data eventually some darts would land outside the 20, 5, and 1 segments and allow for the identification of the correlation. But with professional darts players, we would be waiting a very long time for that to happen. is not the cause of this problem. Moreover it is easy to see that this problem also applies to double regions.
@@
Of course, this correlation identification issue does not arise with the Basic and (in the case of the double regions) IJOC benchmark models. In the Basic and IJOC models, we fit a single bivariate-normal skill model to all the (doubles and trebles) data and all the doubles data, respectively. Correlation identification is therefore feasible for these models but unfortunately, as we saw in Section 6, these models do not fit the data well and are dominated by the DM-based models. In fact, we believe that correlation is likely to be close to zero for all players and target regions and evidence for this is supplied by the one target region - DB - where the correlation coefficient can be identified. Indeed we can see from the fitted BSPN models in Figure 2 that the correlations seem reasonably close to zero and this applies quite generally to all players as may be seen from Figure 9 in Appendix D.
In summary, it is perhaps best to view the bivariate-normal skill model as a mechanism that allows us to (i) fit the raw- or pseudo-counts data and (ii) then target any precise target on the dartboard. In Table 2, we display the observed versus fitted scoring percentages for four players when the bivariate-normal is fit to the raw data when T20 is the target. This ability to fit the observed data is not surprising as the bivariate normal has five free parameters available to fit just six observed outcomes.
@@ 8.1 The Impact of Using Misspecified Models in ZSGs
Anderson 0.8 3.5 0.4 1.8 0 0.1 0 0.1
Aspinall 0.3 1.8 1.1 4.9 0 0.2 0.1 0.3
Chisnall 2.1 9.3 0.4 1.7 0 0.2 0 0.2
Clayton 2.8 12.1 2.1 9.7 0 0.1 0 0.2
Cross 1 4.8 1.2 5.1 0 0.2 0 0.1
Cullen 1.5 6.8 0.8 3.4 0 0.1 0 0.2
@@
Gurney 0.4 1.9 0.9 3.5 0 0.2 0 0.2
Lewis 1 4.7 0.2 0.9 0 0.2 0 0.2
Price 0.7 3.3 0.5 2.5 0 0.2 0 0.2
@@ 8.2 A Single Target Versus Multiple Targets Per Target Region
than a single-action set
@@ 9 Conclusions and Further Research
We performed an exploratory data analysis on a data-set for the top 16 professional players from the 2019 season. We identified several problems that arose due to natural limitations in the data and proposed an empirical Bayesian approach based on the DM distribution to overcome some of these problems. Specifically we introduce two DM-based skill models where the first model borrows strength from other darts players and the second model borrows strength from other regions of the dartboard. We evaluate these models through the use of proper scoring rules and find that they outperform other benchmark models in terms of both statistical and practical significance. We also used our DM-models to estimate the importance of being able to target every square millimeter of the dartboard, and also to analyze specific situations that arose in real-world darts matches during the 2019 season.
@@
(Authors’ names blinded for peer review)
@@ D.2 Statistical Significance of Scoring Results
We also conducted a series paired t-tests to test whether the difference in test scores were significant different from zero. We did this for each proper scoring rule (Brier or Spherical) and for the trebles, doubles and DB regions. For example, referring to the results in Table 6 we could compare the Brier Score aggregated across the treble regions for the BSP skill model versus the IJOC skill model by running a paired t-test with 16 observations. Because each player borrows strength from other players in the BSP, BSPN and BSPN skill models, the samples for the t-tests involving these skill models are not strictly independent. Nonetheless, the dependence is likely to be very weak owing to the fact that the test sets for all the proper scoring rule estimates were independent across players. We report our results in Tables 11 and 12 for the Brier and Spherical scores, respectively.

### v3

@@ 7 Visualizing the Fitted DM Models
In Section 6 we saw that the DM-based models outperform the benchmark, i.e. non-DM, models in terms of proper scoring rules. In this section we will demonstrate this visually by comparing the BSPN model with an analogous model that is fitted using the raw data. In particular, recall that the BSPN model uses a separate skill model (5) for each double and treble region and fits each of these models using the pseudo-fractions $(\widehat{\alpha}_{k}+x_{jk})/\sum_{l=1}^{K}(\widehat{\alpha}_{l}+x_{jl})$. We will compare the BSPN model with the same model but with all the $\widehat{\alpha}_{k}$’s set to zero so that only the raw fractions are used. We will refer to this latter model as the raw-normal (RN) model.
@@
Severe problems can occur when fitting the RN model to double regions, including even D20, which accounts for over 25% of all doubles in the data-set. CEs for D20 and D14 are displayed in Figure 3. In the case of D20, for example, while all 4 players have observations in D20, S20, and M, these are the only regions covered by Cross, while Cullen (S1), van Gerwen (D5), and Gurney (D5) only have a coverage of 4 when targeting D20. With such limited coverage the likelihood function (which the EM algorithm seeks to optimize) will be quite flat in large neighborhoods and terminating at a reasonable (local) maximum is quite unlikely although it can happen. For example, the RN CEs for van Gerwen and Gurney on D20 look reasonable despite only having a coverage of 4. The quality of fits deteriorates as we consider doubles with even less data. There were only 312 darts targeted at D14, for example, which averages out to approximately 19.5 darts per player. It is no surprise then that the D14 coverages are low for all players including Cross (4), Cullen (3), van Gerwen (4), and Gurney (3) whose fits are displayed in the lower row of Figure 3. Matters are even worse for the odd doubles with D19 (the best of them) having only 154 darts targeting it for an average of 9.6 darts per player and all but 1 player having a coverage of just 2 or 3.
@@ 7.1 An Aside on Correlation Identification
The superior fits provided by the BSPN models bring to light an interesting issue. We see from Figures 3 and 4 that the main axis of the CEs (for the BSPN fits) tends to be aligned with the principal direction of the target region in question. For example, in the case of D14 we see the CEs tend to be oriented in a north-south direction which is also the general orientation of the D14 bed. Similarly, we see the CEs for D20 are oriented east-west which is also the orientation for the D20 bed. Indeed, as may be seen from Figures 10-14 in Appendix D.3, this behavior persists more generally. This is a natural consequence of the bivariate-normal skill model as it stretches the variance along the main axis of the target area in order to fit the fractions of darts that landed in the neighboring doubles / trebles regions. A bivariate distribution with fatter tails, e.g. the bivariate $t$ distribution, might better fit this problem but owing to space considerations as well as other issues related to correlation identification (see below) and interpreting ${\bm{\mu}}$ (see Appendix C.1) we leave this for future research.
Considering this tendency to stretch the variance along the main axis of the target area, we believe the resulting (non-zero) correlations are spurious due to a problem intrinsic to the data-set. In particular, we do not know the exact $(x,y)$-coordinates of where a dart lands. Instead we only know the region, e.g. S20, T20, etc., and this means identifying the correlation parameter is very difficult. To see this, consider Figure 5 which displays three synthetic data-sets each consisting of 100 darts aimed at T20. It is clear the skill-models generating these data-sets have very different correlations with the first, second, and third data-sets displaying positive, negative, and (approximately) zero correlations, respectively. However, all three data-sets have identical outcomes, i.e. 37, 54, 6, 1, and 2 darts hitting T20, S20, S5, T1, and S1, respectively. As such, the three data-sets would have identical likelihood functions and there is no way for us to identify the correct sign (never mind value) of the correlation. While we have no darts hitting T5 in these artificial data-sets, it should be clear that not having full coverage [^16] It is true, however, that in the limit of infinite data eventually some darts would land outside the 20, 5, and 1 segments and allow for the identification of the correlation. But with professional darts players, we would be waiting a very long time for that to happen. is not the cause of this problem. Moreover it is easy to see that this problem also applies to double regions.
Note: The blue “+” point is the intended target, and the red “$\times$” points are the landing locations. The three panels have the same numbers of hits in each region: 37 in T20, 54 in S20, 6 in S5, 1 in T1, and 2 in S1.
@@
Of course, this correlation identification issue does not arise with the Basic and (in the case of the double regions) IJOC benchmark models. In the Basic and IJOC models, we fit a single bivariate-normal skill model to all the (doubles and trebles) data and all the doubles data, respectively. Correlation identification is therefore feasible for these models but unfortunately, as we saw in Section 6, these models do not fit the data well and are dominated by the DM-based models.
In summary, it is perhaps best to view the bivariate-normal skill model as a mechanism that allows us to (i) fit the raw- or pseudo-counts data and (ii) then target any precise target (e.g. every square-millimeter) on the dartboard. In Table 2, we display the observed versus fitted scoring percentages for four players when the bivariate-normal fits to the raw data when T20 is the target. This ability to fit the observed data is not surprising as the bivariate normal has five free parameters available to fit just six observed outcomes. While we could enforce zero correlations within the EM algorithm, we found this resulted (in unreported results) in a deterioration of the skill model as measured by proper scoring rules. As such, rather than attach any physical meaning to the fitted correlation parameters, we simply view them as additional degrees of freedom which allows us to improve the forecasting capabilities of the skill models.
@@ 8.1 The Impact of Using Misspecified Models in ZSGs
Anderson 0.8 3.5 0.4 1.8 0.0 0.1 0.0 0.1
Aspinall 0.3 1.8 1.1 4.9 0.0 0.2 0.1 0.3
Chisnall 2.1 9.3 0.4 1.7 0.0 0.2 0.0 0.2
Clayton 2.8 12.1 2.1 9.7 0.0 0.1 0.0 0.2
Cross 1.0 4.8 1.2 5.1 0.0 0.2 0.0 0.1
Cullen 1.5 6.8 0.8 3.4 0.0 0.1 0.0 0.2
@@
Gurney 0.4 1.9 0.9 3.5 0.0 0.2 0.0 0.2
Lewis 1.0 4.7 0.2 0.9 0.0 0.2 0.0 0.2
Price 0.7 3.3 0.5 2.5 0.0 0.2 0.0 0.2
@@ 8.2 A Single Target Versus Multiple Targets Per Target Region
rather than a single-action set
@@ 9 Conclusions and Further Research
We performed an exploratory data analysis on a data-set for the top 16 professional players from the 2019 season. We identified several problems that arose due to natural limitations in the data and proposed an empirical Bayesian approach based on the DM distribution to overcome some of these problems. Specifically we introduce two DM-based skill models where the first model borrows strength from other darts players and the second model borrows strength from other regions of the dartboard. We evaluate these models through the use of proper scoring rules and find that they outperform other benchmark models in terms of both statistical and practical significance. We also used our DM-based models to estimate the importance of being able to target every square millimeter of the dartboard, and also to analyze specific situations that arose in real-world darts matches during the 2019 season.
@@
Acknowledgements
We are very grateful to Christopher Kempf (Twitter @ochepedia), Statistical Analyst for the PDC for providing us with the data underlying this work and for some very insightful conversations. All errors are our own.
@@ Acknowledgements
Martin B. Haugh
Department of Analytics, Marketing & Operations
Imperial College Business School, Imperial College
m.haugh@imperial.ac.uk
Chun Wang
Department of Management Science and Engineering
School of Economics and Management, Tsinghua University
wangchun@sem.tsinghua.edu.cn
@@ D.2 Statistical Significance of Scoring Results
We also conducted a series of paired t-tests to test whether the differences in test scores were significantly different from zero. We did this for each proper scoring rule (Brier or Spherical) and for the trebles, doubles and DB regions. For example, referring to the results in Table 6 we could compare the Brier Score aggregated across the treble regions for the BSP skill model versus the IJOC skill model by running a paired t-test with 16 observations. Because each player borrows strength from other players in the BSP, BSPN and BSPN skill models, the samples for the t-tests involving these skill models are not strictly independent. Nonetheless, the dependence is likely to be very weak owing to the fact that the test sets for all the proper scoring rule estimates were independent across players. We report our results in Tables 11 and 12 for the Brier and Spherical scores, respectively.